The length of the longer leg of a right triangle is
6m more than twice the length of the shorter leg. The length of the hypotenuse is 9m more than twice the length of the shorter leg. Find the side lengths of the triangle. Thank you

Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



The sides are , , and . Plug these expressions into the Pythagorean Theorem and solve for . Then calculate and .



Note that any negative value for would be absurd; exclude any negative root.

John

My calculator said it, I believe it, that settles it

Answer by stanbon(75887)

The length of the longer leg of a right triangle is
6m more than twice the length of the shorter leg. The length of the hypotenuse is 9m more than twice the length of the shorter leg. Find the side lengths of the triangle. Thank you
-------
shorter leg: x
longer leg:: 2x+6
hypot:: 2x+9
-----------------------
Equation:
(2x+9)^2 = x^2 + (2x+6)^2
------
4x^2 + 36x + 81 = x^2 + 4x^2 + 24x + 36
-------
x^2 -12x - 45 = 0
------
x = 15
----
Ans:
shorter leg = 15
longer leg = 36
----------
Cheers,
Stan H.
---------------